[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Oct 31, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Oct 31 13:23:25 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
Several TAFB products were not being sent or posting on the web yesterday  
afternoon and evening. These were caused a problem on NHC's web server and  
on the ldm on the logger server that picks up products after they are sent  
out through AWIPS. Dave switched to the backup web server yesterday  
afternoon, and Matt Sardi repaired the ldm late yesterday evening. Dave is  
working with NCO today to fix the primary NHC web server.
HSU:
There were some problems yesterday with the TWO products not updating.  
These were caused by the same two issues mentioned above for TAFB.
Other items:
There was an AC problem on Sunday in the computer room Sunday night that  
required the WFO to shut down their AWIPS. The temperature reached 95 F at  
one point. It is possible that the high temperatures were related to  
various NHC data flow problems mentioned above. Mark will coordinate with  
Jann to see if we can implement a system to better monitor the computer  
room so we can take action before the computer room gets that hot.

NHC RFCs:
RFCs 206 and 207 will be implement today to add scripts that will allow  
forecasters to remove the storm surge watch warning and inundation graphic  
from the NHC web page at the end of a TC. Previously, only the NHC  
webmaster could do that.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161031/2b83183a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list